Kyky232:

If segment FD = 30 units and segment FE = 55 units, what is the value of a? Round the solution to the nearest hundredth. (pic in replies) 22.62 28.44 41.41 61.39

Kyky232:

1 attachment
KaityK24:

which one do u think it is

simsharrison:

ok, so F is a 90* angle and the measurements of a triangle =180*

snowflake0531:

because you're wanting to know the angle measure of a, you know the opposite side and the adjacent side, so, which of sine, cosine, and tangent would you use

Kyky232:

i dont know.... im bad with these...

snowflake0531:

do you know soh cah toa

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 do you know soh cah toa \(\color{#0cbb34}{\text{End of Quote}}\) no..

snowflake0531:

well sin(x) = opposite/hypotenuse cos(x) = adjacent/hypotenuse tan(x) = opposite/adjacent and becuase we want opposite/adjacent, which one are we going to use

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 well sin(x) = opposite/hypotenuse cos(x) = adjacent/hypotenuse tan(x) = opposite/adjacent and becuase we want opposite/adjacent, which one are we going to use \(\color{#0cbb34}{\text{End of Quote}}\) tan(x)?

snowflake0531:

yes

snowflake0531:

so, like how i wrote it what would the equation be like

simsharrison:

Ok, what you want to do issss: start with what we know: angle measure of F = 90* right angel right? So since we are finding the angle measure of A, we will have to use one of those formulas that @snowflake0531 is telling you about.

Kyky232:

tan(a)?

snowflake0531:

yes! so tan(a) = what

snowflake0531:

remember, tan(a) = opposite/hypotenuse so, what is it

snowflake0531:

anyways lol tan(a) = 30/55 so, i recommend putting tan^-1 (30/55) into your calc

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 anyways lol tan(a) = 30/55 so, i recommend putting tan^-1 (30/55) into your calc \(\color{#0cbb34}{\text{End of Quote}}\) good job - congrats !

Kyky232:

tan(A)=30/55

Kyky232:

ope so it sent late-

snowflake0531:

so now put it into mathway or something tan(x) = 30/55 and figure out what the angle is

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 so now put it into mathway or something tan(x) = 30/55 and figure out what the angle is \(\color{#0cbb34}{\text{End of Quote}}\) it says .6?

Kyky232:

im so confused...

snowflake0531:

Kyky232:

oh i used the wrong calculator. mathway is blocked-

AddyM:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Kyky232 oh i used the wrong calculator. mathway is blocked- \(\color{#0cbb34}{\text{End of Quote}}\) oof everything is blocked o nour laptops i swear.

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AddyM \(\color{#0cbb34}{\text{Originally Posted by}}\) @Kyky232 oh i used the wrong calculator. mathway is blocked- \(\color{#0cbb34}{\text{End of Quote}}\) oof everything is blocked o nour laptops i swear. \(\color{#0cbb34}{\text{End of Quote}}\) fr tho

snowflake0531:

hmmm, same lol, school blocks so much haha

Kyky232:

wait so is it B?

snowflake0531:

i think